In a certain experiment, the probability of an error is 1%, or 0.01. If the experiment is repeated four times, find the probability that at least one of the trials will result in an error
Answer by stanbon(75887)

In a certain experiment, the probability of an error is 1%, or 0.01. If the experiment is repeated four times, find the probability that at least one of the trials will result in an error
----
P(error) = 0.01
P(no error) = 0.99
------
Binomial Problem with n = 4
---
P(at least one error) = 1 - P(no errors in 4 trials)
---
= 1 - (0.99)^4
---
= 1 - 0.9606
---
= 0.0394
===============
Cheers,
Stan H.
============